# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ Appendix B EFFECTIVE MODULUS OF LER
Figure 31: (a) and (b) show the relation between the average neighbor-keeping ratio among slow particles, denoted as $\bar{C}$, and the ratio of particles that have experienced cage jump in the system, denoted as $\phi$, for sample A and sample B, respectively, at various $\dot{\gamma}$. (c) and (d) show the results of $w$ (symbols), found from Eq. 38, as a function of $\dot{\gamma}$ for sample A and sample B, respectively. In (c) and (d), horizontal dashed lines denote the mean values of $w$ averaged over all $\dot{\gamma}$ for sample A and sample B.
@@
The values of $w$ obtained by the preceding equation for sample A and sample B are given in Fig. 31 (c) and (d), respectively.
@@ Appendix C IMPORTANCE OF STRUCTURAL FEATURES
Table 1: The $15$ important features for two samples.
sample A sample B
1 $2$ 00 00 $1.3$ $2$ 00 00 $1.3$
2 $2$ $2$ $-2$ $0.9$ $2$ $2$ $-2$ $1.2$
3 $2$ 00 00 $1.1$ $2$ 00 00 $1.4$
4 $2$ 00 00 $1.2$ $2$ 00 00 $0.9$
5 $2$ 00 00 $3.0$ $2$ 00 00 $1.2$
6 00 00 00 $1.0$ $1$ $2$ $-2$ $1.2$
7 $1$ $2$ $-2$ $0.9$ $1$ 00 00 $1.4$
8 $2$ 00 00 $0.8$ $2$ 00 00 $1.1$
9 $2$ 00 00 $1.9$ $2$ 00 00 $2.3$
10 $2$ 00 00 $1.4$ 00 00 00 $1.3$
11 $2$ 00 00 $0.7$ $2$ 00 00 $2.1$
12 $2$ 00 00 $1.1$ $2$ 00 00 $3.4$
13 $2$ $2$ $-2$ $2.9$ $2$ 00 00 $1.5$
14 $2$ 00 00 $1.7$ $2$ 00 00 $0.8$
15 $2$ 00 00 $1.6$ 00 $2$ $-2$ $5.0$

### v2

@@ Appendix B EFFECTIVE MODULUS OF LER
The values of $w$ obtained by the preceding equation for sample A and sample B are given in figure 31 (c) and (d), respectively.
@@ Appendix C IMPORTANCE OF STRUCTURAL FEATURES
Table 1: The $15$ important features for two samples.
sample A sample B
1 22 0 0 $1.3$ 22 0 0 $1.3$
2 22 22 2-2 $0.9$ 22 22 2-2 $1.2$
3 22 0 0 $1.1$ 22 0 0 $1.4$
4 22 0 0 $1.2$ 22 0 0 $0.9$
5 22 0 0 $3.0$ 22 0 0 $1.2$
6 0 0 0 $1.0$ 11 22 2-2 $1.2$
7 11 22 2-2 $0.9$ 11 0 0 $1.4$
8 22 0 0 $0.8$ 22 0 0 $1.1$
9 22 0 0 $1.9$ 22 0 0 $2.3$
10 22 0 0 $1.4$ 0 0 0 $1.3$
11 22 0 0 $0.7$ 22 0 0 $2.1$
12 22 0 0 $1.1$ 22 0 0 $3.4$
13 22 22 2-2 $2.9$ 22 0 0 $1.5$
14 22 0 0 $1.7$ 22 0 0 $0.8$
15 22 0 0 $1.6$ 0 22 2-2 $5.0$